OpenStudy (anonymous):

how do you solve this ? x/2-1=x/5+2

OpenStudy (lalaly):

\[\frac{x}{2}-\frac{x}{5}=2+1\]\[(\frac{x}{2} \times \frac{5}{5})-(\frac{x}{5} \times \frac{2}{2})=3\]\[\frac{5x}{10}-\frac{2x}{10}=3\]\[\frac{3x}{10}=3\]\[3x=10 \times 3\]\[3x=30\]\[x=10\]

OpenStudy (anonymous):

thanks:)

OpenStudy (lalaly):

yaaaaaaaaaaaaaay u gave me my 2000 medal xD

OpenStudy (lalaly):

thankyou hehe

OpenStudy (anonymous):

you gave me my first medal :) thanks im guessing the more medals the better huh

OpenStudy (anonymous):

Hi everyone, im very new at this wedsite; like 5 minutes new so i was wondering how hard the questions could be? Like Xx50=5000+45-S=.... <<<What do u think the answer is??

OpenStudy (anonymous):

hi im new to this too :) so i really wouldent know how to answer:)